STATEMENT OF COMPREHENSIVE INCOME

  • Always apply the income tax expense as 25% of operating profit.

Gross Sales = ₱ 3,822,300

Sales Returns= ₱ 219,700

Gross Purchases = ₱ 1,000,000

Purchase Discounts = ₱ 180,000

Beginning Inventories = ₱ 920,500

Ending Inventories = ₱ 450,800

Distribution Expenses = ₱ 760,200

Administrative Expenses = ₱ 440,900

Other Comprehensive Income = ₱ 210,000


Look for:

A. NET PROFIT MARGIN

B. GROSS PROFIT MARGIN

C. OPERATING PROFIT MARGIN

1
Expert's answer
2021-10-25T09:17:59-0400

COGS= 920500+1000000-180000-450800 =1289700

Sales = 3,822,300- 219,700 = 3602600

Gross Profit = 3602600-1289700 =2312900

Expenses = 760,200+440,900+210,000= 1411100

Operating profit = 2312900-1411100=901800

Income tax = 25/100 x901800= 225450

Net profit = 901800-225450= 676350

1.NET PROFIT MARGIN = (Net profit/revenue) x 100

= (676350/3602600) x 100 =18.8 %

2.GROSS PROFIT MARGIN = (Gross profit/revenue) x100

=(2312900/3602600) x 100= 64.2 %

3.OPERATING PROFIT MARGIN= (Operating profit/revenue) x100

=(901800/3602600)x 100 =25.03%